1.A large tank contains water having a depth of 1.5 m. lf the tank is on an elevator that is descending, determine the horizontal speed of a wave created on its surface if the rate of descent is (a) constant at 6 m/s. (b) accelerated at 1.5 m/s2. (c) accelerated at 9.81 m/s2.


2.. Water flows at 840(103) liters/min through a rectangular channel having a width of 3 m. If yA = 1 m, show that yB = 1.25 m is possible. What is the necessary increase in elevation h of the channel bed to create this depth?



1
Expert's answer
2022-01-20T10:08:32-0500

1.

a)

"v=\sqrt{gh}=3.8~\frac ms,"

b)

"v=\sqrt{(g-a)h}=3.5~\frac ms,"

c)

"v=\sqrt{(g-a)h}=0~\frac ms."

2.

"h=\frac Qg\sqrt{y^2_B-y^2_A}=0.4~m."